# EVolve: a Value-Added Services Platform for Electric Vehicle Charging Stations

### v1

@@ Abstract
A notable challenge in Electric Vehicle (EV) charging is the time required to fully charge the battery, which can range from 15 minutes to 2-3 hours. This idle period, however, presents an opportunity to offer time-consuming or data-intensive services such as vehicular software updates. ISO 15118 referred to the concept of Value-Added Services (VAS) in the charging scenario, but it remained underexplored in the literature. Our paper addresses this gap by proposing EVolve, the first EV charger compute architecture that supports secure on-charger universal applications with upstream and downstream communication. The architecture covers the end-to-end hardware/software stack, including standard API for vehicles and IT infrastructure. We demonstrate the feasibility and advantages of EVolve by employing and evaluating three suggested value-added services: vehicular software updates, security information and event management (SIEM), and secure payments. The results demonstrate significant reductions in bandwidth utilization and latency, as well as high throughput, which supports this novel concept and suggests a promising business model for Electric Vehicle charging station operation.
Index Terms:
Electric Vehicle (EV) Charging, Value-Added Services (VAS), Secure Payments, Latency Reduction, EV Station Data Operation.
@@ I Introduction
With the increasing adoption of Electric Vehicles (EVs), the demand for advanced, secure, and scalable charging infrastructure has become essential [1][2]. Governments and industries worldwide are heavily investing in EV technologies to reduce carbon footprints and transition towards sustainable mobility solutions [3] [4]. However, EVs’ prominence brings its own challenges, particularly related to charging infrastructure, security, and user experience.
EV chargers are becoming increasingly ubiquitous, yet their exposure to potential threats necessitates robust security measures. Standards such as ISO 15118 [5] [6] [7] have been introduced to address these security and communication challenges. Still, significant gaps remain, especially in supporting data-rich and secure interactions between vehicles, chargers, and cloud-based services [8] [2]. Furthermore, the charging process is time-consuming, often leaving users idle during the charging period. This idleness allows users to engage in relevant activities, potentially transforming the charging experience.
@@
This work introduces EVolve, a platform designed to elevate EV chargers from energy dispensers to robust Edge Computing (EC) devices. EVolve brings a novel concept transforming the EV charging ecosystem by supporting Value Added Services (VASs) in the charging environment. By enhancing data communication and security and introducing innovative services, EVolve addresses critical gaps in the existing infrastructure.
Leveraging Vehicular Edge Computing (VEC), EVolve expands the functionality of EV chargers, enabling secure, data-driven services that optimize user experience, improve grid utilization, and safeguard data confidentiality. The platform supports applications like vehicular software updates, security detection and response, e-payments, grid load balancing, and driving UX enhancements, thereby bridging the gap between Operational Technology (OT) and Information Technology (IT), while also being compliant to the ISO 15118 standard.
The empirical evaluation of EVolve demonstrates its feasibility by implementing and evaluating advanced VASs such as vehicular software updates, Security Information and Event Management (SIEM), and secure payments. The emulation results, mainly latency and throughput, show that the EVolve platform consistently outperforms cellular-based communications like 4G and 5G in most tested scenarios, thanks to the wired Power Line Communication (PLC)/Ethernet connectivity to the EV charger. Remarkably, EVolve outperforms 5G nearly seven times in high-frequency applications like secure micro-payments.
The remainder of this paper is organized as follows: Section 2 reviews the related work that inspired and informed our research. In Section 3, we explore the challenges and opportunities in the domain and provide a detailed overview of our proposed architecture. Section 4 focuses on the implementation and evaluation of three Value-Added Services, detailing how the identified challenges were addressed. This section also includes an analysis of the evaluation results, along with considerations of system overhead and requirements. Finally, Section 5 concludes the paper, summarizing key findings.
@@ II Related Works
The ISO 15118 (first published in 2014) serves as the foundational standard for enabling high-level communication between EVs and Supply Equipment Communication Controller (SECC). It also briefly highlights the importance of having Value-added serivces (VAS) in the future, that leverage the edge-like data and compute capabilities of EV charger. Nevertheless, this topic has not been addressed deeply in literature, and no generic architectures have been proposed, unless primitive use-cases like online payments.
@@
VEC brings computational and storage resources closer to vehicles, addressing limitations in cloud-dependent systems by reducing latency and enhancing real-time processing. As highlighted in [10] and [11], edge computing enables a wide range of applications, including software updates [12], map download, and predictive maintenance [13], all of which are essential for modern EV infrastructure.
The integration of VAS into the EV charging ecosystem has been driven by efforts to utilize idle charging time more effectively. One prominent example is the Verband Deutscher Verkehrsunternehmen (VDV) resolution 261 [14], which introduced a VAS to precondition electric buses during charging. This service, compliant with ISO 15118, ensures that bus interiors are climatized, enhancing passenger comfort and operational efficiency while conserving battery power for driving. Despite initial implementation challenges, its expansion to countries like the Netherlands demonstrates the feasibility of such VAS in public transportation and sets a benchmark for future developments [8] [15]. Our work takes the further steps through presenting a generic architecture to enable several applications.
@@
Since its introduction in 2013 [16], VAS has evolved to expand the value proposition of EV charging stations. Software updates represent one of the earliest practical applications, addressing the underutilization of EV chargers as service delivery points. By leveraging the proximity of chargers to vehicles, works such as ScalOTA [12] have developed secure, end-to-end update protocols, utilizing idle charging time to ensure efficient and secure firmware updates for Electronic Control Units (ECUs). This approach reduces reliance on workshops and enhances convenience for both Original Equipment Manufacturers (OEMs) and users, highlighting the role of VEC in modernizing the EV ecosystem. Furthermore, the industry sees edge computing as an essential component for the delivery of comfort and convenience applications [2], further emphasizing the importance of integrating VAS into the EV infrastructure. EVolve generalizes the OTA update application of ScalOTA to cover more applications and support OEM-agnostic vehicles.
@@
While ISO 15118 provides a solid foundation for secure EV communication, its gaps in VAS practical implementation and scalability highlight the need for complementary solutions. Edge computing addresses these limitations by providing a distributed data processing architecture that supports diverse applications with enhanced security and efficiency. By leveraging edge-based platforms like Edgex [17] and Baetyl [18], which integrate microservices and advanced communication protocols, it is possible to deploy secure, scalable, and user-centric VAS in the EV ecosystem. EVolve’s implementation benefits greatly from the integration of these platforms to facilitate building a solid end-to-end solution.
@@ III-A Overview
EVolve is a VEC-based platform designed to bridge the gap left by ISO 15118 in defining and supporting VASs within the EV charging ecosystem. By leveraging EV chargers’ communication and computation capabilities, the platform enables VASs to create a secure, robust environment for running diverse services. EVolve also follows modular architecture and standard APIs and protocols to ease its integration and facilitate building new applications. EVolve emphasizes both security and performance, addressing the lack of a systematic approach to functional, performance, and security requirements for VASs in the literature.
@@ III-B Challenges and Opportunities
EV users suffer from battery “range anxiety”, which slows down the EVs adoption. While there are huge efforts to reduce the charging time and optimize battery utilization, extending the EV chargers with data, computing, and communication capabilities entails huge opportunities. First, it reduces the psychological feeling of “time waste”. Second, it improves the utilization of cellular infrastructure resources (bandwidth and Mobile Edge Compute (MEC)), interference, and reduces the latency of services. Third, automotive applications can now benefit from the diverse data sharing of several fleets and models at the EV charging station. Fourth, it enables a wide set of new services in the connected automotive ecosystem, which was impossible with the constrained vehicle capabilities or overwhelming the IT infrastructure. This opens new business opportunities where EV charging operators can support automotive applications through offloading heavyweight computing or communication to the EV chargers or downloading huge amounts of data (e.g., maps, software updates, etc.).
@@ III-C2 Layer 2 – Microservices and Libraries
The second layer hosts essential microservices and libraries that ensure the platform meets its cybersecurity requirements while enhancing the performance of VASs. Key modules include:
  • Security: Cryptographic libraries, protocols, and secure hardware module drivers to support secure operations, storage, and sensitive data computation.
  • Charging Stack: Ensures reliable charging operations, fulfilling critical high-availability needs.
  • AI and Machine Learning: AI processing software libraries to support offloaded data analytics and inference, enabling advanced VAS tasks while maintaining efficiency.
  • Cache Storage: Provides efficient caching to reduce latency and address availability concerns.
  • Telemetry & Communication: Collects data and ensures safe export to the cloud through a secure connection.
  • Resilience: Delivers fault tolerance and recovery through monitoring, redundancy, and proactive failure handling to ensure system reliability.
@@ III-C4 Layer 3 – Value-Added Services (VASs)
The top layer builds on the capabilities of the lower layers to deliver advanced functionalities such as vehicular software updates, maps, entertainment, user experience enhancements, security information and event management (SIEM), and secure payments. These services leverage the underlying architecture to meet the platform’s usability and cybersecurity requirements, ensuring scalability and adaptability. We considered employing three VASs with different requirements:
  • Vehicular Software Updates: Over-the-Air firmware/software updates.
  • SIEM: Integrates telemetry and analytics to monitor, identify, and mitigate threats.
  • Secure Payments: Payment support for charging and VASs, for untrusted stakeholders.
@@ III-C5 Application Programming Interface (API)
Certain VASs may have complementary components or require heavier processing capabilities hosted on the cloud. Thus, seamless communication between the EV charger and the cloud is essential to enable these functionalities. To maintain data security, all communication between the charger and the cloud is encrypted, ensuring the integrity and confidentiality of transmitted information. Integration with the cloud leverages widely adopted internet protocols, such as HTTPS and MQTTS, providing flexibility and a developer-friendly environment for data export and processing. This communication model supports:
  • Offloading computationally intensive tasks from the charger to the cloud.
  • Synchronization of logs and data analytics for advanced VASs, such as predictive maintenance or grid load balancing.
  • Secure data export for third-party applications or services.
The cohesive integration of vehicle-side and cloud-side APIs enables the platform to serve as a robust and flexible intermediary, bridging the gap between edge and centralized systems while maintaining the core principles of security, performance, and scalability.
@@ IV-B Experimental Setting
Configurations Connection TPS (Mbps) Latency (ms) PLR
@@
TABLE I: Network Performance Comparison for Different Connections
Following the measurements, the collected points were plotted in a time series manner for visualization. This first analysis, shown in Fig. 2, illustrates the stability and faster performance of wired connections (EVolve10, EVolve100, EVolve1G).
@@ IV-C1 Vehicular Software Updates
Fig. 3 shows the flow for the software updates VAS. This service uses a publish/subscribe (pub/sub) mechanism to receive notifications of available firmware updates from an Image Repository (IR). The EV charger acts as the Update Data Broker (UDB), downloading updates securely from the cloud if they are not cached, authenticating them, and implementing version control and rollback mechanisms. This architecture minimizes network reliance and ensures that updates are securely delivered and applied without compromising the system.
@@ IV-C2 SIEM
For the SIEM VAS, which manages both large data logs and smaller FL parameters, we emulated communication-based on a CAN dataset, representing 10 minutes of data totaling approximately 85 MB [24]. In this test, the client sent a 100 MB log file and received a simple ACK from the server, emulating the log upload to the SIEM service. We also tested a small download of 1 MB to evaluate downstream performance in receiving FL parameter updates. However, this case had the same result as the software update. Thus, we chose to omit it.
@@ IV-C3 Secure Payment
Figure 6: Latency for Software Updates.
Figure 7: Latency of SIEM Operations.
@@ V Conclusion
The EVolve platform demonstrates the potential to transform EV chargers into secure, efficient, and scalable hubs for VASs. Evaluation results highlight its ability to leverage edge computing and advanced communication technologies, with the emerging 100 Mbps PLC standard surpassing 4G and 5G in stability and throughput, making it ideal for latency-sensitive tasks like vehicle-to-charger communication.
Secure payment VAS benefits significantly from PLC’s low latency and high bandwidth, providing a secure and efficient alternative to traditional methods. Vehicular software updates and SIEM applications also show marked improvements over 4G, even in a 10 Mbps PLC environment, underscoring the platform’s scalability and adaptability.
These results confirm EVolve as a robust solution for advancing smart EV charging systems, bridging gaps in ISO 15118, and enabling secure, scalable VAS deployment while balancing edge and cloud resources for optimal performance.

### v2

@@ Abstract
A notable challenge in Electric Vehicle (EV) charging is the time required to fully charge the battery, which can range from 15 minutes to 2-3 hours. However, this idle period for the EV presents an opportunity to offer time-consuming or data-intensive services such as vehicular software updates. ISO 15118 referred to the concept of Value-Added Services (VASs) in the charging scenario, but it remained underexplored in the literature. Our paper addresses this gap by proposing EVolve, the first EV charger compute architecture that supports secure on-charger universal applications with upstream and downstream communication. The architecture covers the end-to-end hardware/software stack, including standard API for vehicles and IT infrastructure. We demonstrate the feasibility and advantages of EVolve by employing and evaluating three suggested value-added services: vehicular software updates, security information and event management (SIEM), and secure payments. The results demonstrate significant reductions in bandwidth utilization and latency, as well as high throughput, which supports this novel concept and suggests a promising business model for Electric Vehicle charging station operation.
@@ I Introduction
With the increasing adoption of EVs, the demand for advanced, secure, and scalable charging infrastructure has become essential [15][2]. Governments and industries worldwide are heavily investing in EV technologies to reduce carbon footprints and transition towards sustainable mobility solutions [5] [27]. However, EVs’ prominence brings its own challenges, particularly related to cybersecurity risks in both vehicles and charging systems, limited and inconsistent charging infrastructure, and frustrating user experience.
Standards such as ISO 15118 [21] [22] [23] have been introduced to address these security and communication challenges. Still, significant gaps remain, especially in supporting data-rich and secure interactions between vehicles, chargers, and cloud-based services [7] [2]. Additionally, range anxiety and idle charging time remain major issues for EVs, impacting adoption rates. Drivers often worry about reaching a charging station before running out of charge, only to encounter long waits due to occupied or overstayed chargers. This mix of uncertainty and delay makes the overall charging experience frustrating and unreliable. However, enabling productive or engaging activities during this period could shift user perception, reduce concerns, and benefit the industry by enhancing the overall charging experience.
@@
This work introduces EVolve, a platform designed to elevate EV chargers from energy dispensers to robust Vehicular Edge Computing (VEC) devices. EVolve brings a novel concept transforming the EV charging ecosystem by supporting Value Added Services (VASs) in the charging environment. By enhancing data communication and security and introducing innovative services, EVolve addresses critical gaps in the existing infrastructure.
Our contributions are as follows: 1) the EVolve, a novel VEC architecture that transforms EV chargers into secure, data-driven hubs bridging OT and IT, compliant with ISO 15118; 2) the integration of VASs such as software updates, SIEM, and secure payments; 3) the practical implementation and feasibility demonstration of EVolve; and 4) empirical performance evaluation showing significantly lower latency and higher throughput compared to 4G/5G, with up to a seven-fold improvement over 5G in high-frequency tasks like secure payments, enabled by wired Power Line Communication (PLC) as specified by ISO 15118. Moreover, our architecture is the base for our work on Security Orchestration, Automation and Response for EVs on [12], showcasing the platform’s flexibility and practicality.
The remainder of this paper is organized as follows: Section 2 reviews the related work that inspired and informed our research. In Section 3, we explore the challenges and opportunities in the domain and provide a detailed overview of our proposed architecture. Section 4 focuses on the implementation and evaluation of three VASs, detailing how the identified challenges were addressed. This section also includes an analysis of the evaluation results and considerations of system overhead and requirements. Finally, Section 5 concludes the paper, summarizing key findings.
@@ II Related Works
The ISO 15118 (first published in 2014) serves as the foundational standard for enabling high-level communication between EVs and Supply Equipment Communication Controller (SECC). It also briefly highlights the importance of having VASs in the future, that leverage the edge-like data and compute capabilities of EV charger. Nevertheless, this topic has not been addressed deeply in literature, and no generic architectures have been proposed, unless primitive use-cases like online payments.
@@
VEC brings computational and storage resources closer to vehicles, addressing limitations in cloud-dependent systems by reducing latency and enhancing real-time processing. As highlighted in [19] and [4]. Edge computing enables a wide range of applications, including software updates [26], map download, and predictive maintenance [16], all of which are essential for modern EV infrastructure.
Previous studies [25] [28] have explored the role of EVs as both energy prosumers and computing nodes, optimizing charging, computation, and travel costs through coordinated management of EVs, the Power Distribution Network, and the Computing Power Network, a fusion of EV and edge server. While these works primarily focus on optimizing EV-side operations, our research takes a different approach by emphasizing the SECC and edge servers, ensuring that the underlying infrastructure can support such capabilities. Furthermore, we adopt the concept of VASs, facilitating the deployment of these technologies rather than solely optimizing their operation. Consequently, our work complements rather than overlaps with these prior studies.
@@
Since its introduction in 2013 [24], VAS has evolved to expand the value proposition of EV charging stations. Software updates represent one of the earliest practical applications, addressing the underutilization of EV chargers as service delivery points. By leveraging the proximity of chargers to vehicles, works such as ScalOTA [26] have developed secure, end-to-end update protocols, utilizing idle charging time to ensure efficient and secure firmware updates for Electronic Control Units (ECUs). This approach reduces reliance on workshops and enhances convenience for both Original Equipment Manufacturers (OEMs) and users, highlighting the role of VEC in modernizing the EV ecosystem. Furthermore, the industry sees edge computing as an essential component for the delivery of comfort and convenience applications [2], emphasizing the importance of integrating VASs into the EV infrastructure. EVolve generalizes the OTA update application of ScalOTA to cover more applications and support OEM-agnostic vehicles.
The integration of VASs into the EV charging ecosystem has been driven by efforts to utilize idle charging time more effectively. One prominent example is the VDV resolution 261 [6], which introduced a VAS to precondition electric buses during charging in Germany. This service, compliant with ISO 15118, ensures that bus interiors are climatized, enhancing passenger comfort and operational efficiency while conserving battery power for driving.
@@
While ISO 15118 provides a solid foundation for secure EV communication, its gaps in VAS practical implementation and scalability highlight the need for complementary solutions. Edge computing addresses these limitations by providing a distributed data processing architecture that supports diverse applications with enhanced security and efficiency. By leveraging edge-based platforms like Edgex [9] and Baetyl [18], which integrate microservices and advanced communication protocols, it is possible to deploy secure, scalable, and user-centric VASs in the EV ecosystem. EVolve’s implementation benefits greatly from the integration of these platforms to facilitate building a solid end-to-end solution.
@@ III-A Overview
EVolve is a VEC-based platform designed to bridge the gap left by ISO 15118 in defining and supporting VASs within the EV charging ecosystem. By leveraging EV chargers’ communication and computation capabilities, the platform enables VASs to create a secure, robust environment for running diverse services. EVolve also follows modular architecture, standard APIs, and protocols to ease its integration and facilitate building new applications. EVolve emphasizes both security and performance, addressing the lack of a systematic approach to functional, performance, and security requirements for VASs in the literature.
@@ III-B Challenges and Opportunities
Although there are huge efforts to reduce the charging time and optimize battery utilization, extending the EV chargers with data, computing, and communication capabilities entails huge opportunities. First, it reduces the psychological feeling of “time waste”. Second, it improves the utilization of cellular infrastructure resources (bandwidth and Mobile Edge Compute (MEC)), interference, and reduces the latency of services. Third, automotive applications can now benefit from the diverse data sharing of several fleets and models at the EV charging station. Fourth, it enables a wide set of new services in the connected automotive ecosystem, which was impossible with the constrained vehicle capabilities or overwhelming the IT infrastructure. This opens new business opportunities where EV charging operators can support automotive applications through offloading heavyweight computing or communication to the EV chargers or downloading huge amounts of data (e.g., maps, software updates, etc.).
@@ III-C2 Layer 2 – Microservices and Libraries
The second layer incorporates critical microservices and libraries designed to meet stringent cybersecurity requirements and optimize the performance of VASs. This layer encompasses essential modules such as Security, which includes cryptographic libraries, protocols, and secure hardware module drivers for secure operations, storage, and sensitive data computation; Charging Stack, ensuring reliable charging operations with high availability; AI and Machine Learning, featuring AI processing software libraries to support offloaded data analytics and inference, enhancing VAS efficiency; Cache Storage, providing efficient caching to minimize latency and improve availability; Telemetry & Communication, collecting data and ensuring secure cloud export; and Resilience, delivering fault tolerance and recovery through monitoring, redundancy, and proactive failure handling to guarantee system reliability.
@@ III-C4 Layer 3 – Value-Added Services (VASs)
The top layer delivers advanced functionalities, including Vehicular Software Updates for over-the-air firmware/software updates, SIEM integrating telemetry and analytics for threat monitoring and mitigation, and Secure Payments supporting charging and VAS transactions for untrusted stakeholders, alongside maps, entertainment, and user experience enhancements, leveraging the lower layer’s capabilities to meet usability and cybersecurity requirements while ensuring scalability and adaptability.
@@ III-C5 Application Programming Interface (API)
Seamless, encrypted communication between the EV charger and the cloud, utilizing HTTPS and MQTTS, is crucial for VASs requiring cloud-hosted components or heavier processing, supporting: offloading computationally intensive tasks, synchronizing logs and data analytics for advanced VASs like SIEM, predictive maintenance and grid load balancing, and secure data export for third-party applications. This cohesive vehicle-side and cloud-side API integration allows the platform to act as a robust intermediary, bridging edge and centralized systems while upholding security, performance, and scalability.
@@ IV-B Experimental Setting
TABLE I: Network Performance Comparison for Different Connections
Configuration Connection TPS (Mbps) Latency (ms) PLR
@@
Following the measurements, the collected points were plotted in a time series manner for visualization. This first analysis, shown in Fig. 2, illustrates the higher stability of wired connections (EVolve10, EVolve100, EVolve1G).
@@ IV-C1 Vehicular Software Updates
Fig. 3 shows the flow for the software updates VAS. This service uses a pub/sub mechanism to receive notifications of available firmware updates from an Image Repository. The EV charger then downloads updates securely from the cloud if they are not cached, authenticates them, and implements version control and rollback mechanisms. This architecture minimizes network reliance and ensures that updates are securely delivered and applied without compromising the system.
@@ IV-C2 SIEM
Figure 6: Latency of SIEM Operations.
For the SIEM VAS, which manages both large data logs and smaller FL parameters, we emulated communication based on a CAN dataset, representing 10 minutes of data totaling approximately 85 MB [17]. In this test, the client sent a 100 MB log file and received a simple ACK from the server, emulating the log upload to the SIEM service. We also tested a small download of 1 MB to evaluate downstream performance in receiving FL parameter updates. However, this case had the same result as the software update. Thus, we chose to omit it.
@@ V Conclusion
The EVolve platform demonstrates the potential to transform EV chargers into secure, efficient, and scalable hubs for VASs by leveraging edge computing and advanced communication technologies. Evaluation results show that the emerging 100 Mbps PLC standard outperforms 4G and 5G in stability and throughput, making it ideal for latency-sensitive tasks like vehicle-to-charger communication. Secure payment VAS particularly benefits from PLC’s low latency and high bandwidth, providing a secure and efficient alternative to traditional methods. Additionally, vehicular software updates and SIEM applications exhibit significant improvements over 4G, even in a 10 Mbps PLC environment, demonstrating the platform’s scalability and adaptability. These results establish EVolve as a robust solution for advancing smart EV charging systems, addressing gaps in ISO 15118, and enabling secure, scalable VAS deployment while balancing edge and cloud resources for optimal performance.